A bibliography on coordination

Paolo Ciancarini Dipartimento di Scienze dell’Informazione University of Bologna - Italy

               Email: ciancarini@cs.unibo.it

April 9, 2003

This is a list of papers related to coordination models, languages, architectures, and applications.

1 General

In this section we list papers, books, and conference proceedings concerning the general definition of coordination models and languages and related issues [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?].